2 2(10). Determine if the following two equations are exact, and solve it if yes. (a) (? - )dx + (1 - **+ **)dy = 0.

(b) (x + tan- y)dx + **dy = 0.
Answer #1

We have, Equation m dat Nody =0 is exact it am = IN ay Jx 20 c? du + Ct -e2 +24) dy=0 У 24 where ME are -3y N = 1 У X4 42 23(x+y) - dy=0 b (at tanyl dx + 42f1 Here M= rttany xty y²f1 ӘМ by ot y²41 y²+1 IN = 1 and y²41 M = IN dy . Given equation is